OpenStudy (anonymous):

Write an equation of a line in point-slope form that passes through the point (1,2) and whose slope is 3.

OpenStudy (anonymous):

For this, we put it in point-slope. y-y1=m(x-x1). Now all we do is put the info. that we have and make the equation. y-2=3(x-1) y-2=3x-3 Now we add 2 to both sides y=3x-1. This is our final equation.